### v1

# NMPCB: A Lightweight and Safety-Critical Motion Control Framework

Abstract
In multi-obstacle environments, real-time performance and safety in robot motion control have long been challenging issues, as conventional methods often struggle to balance the two. In this paper, we propose a novel motion control framework composed of a Neural network-based path planner and a Model Predictive Control (MPC) controller based on control Barrier function (NMPCB) . The planner predicts the next target point through a lightweight neural network and generates a reference trajectory for the controller. In the design of the controller, we introduce the dual problem of control barrier function (CBF) as the obstacle avoidance constraint, enabling it to ensure robot motion safety while significantly reducing computation time. The controller directly outputs control commands to the robot by tracking the reference trajectory. This framework achieves a balance between real-time performance and safety. We validate the feasibility of the framework through numerical simulations and real-world experiments.
I Introduction
In the design of robot motion planners and controllers, safety has always been the paramount requirement, considering the rapid development of robot systems[1]. The control method based on CBF has garnered significant attention for offering a simple yet safety-guaranteed approach to robot motion control[2]. In particular, the control method that integrates CBF and MPC have demonstrated superior performance in safety control by predicting future states [3]. However, this method often fails to obtain an optimal solution within a short time, especially for nonlinear kinematic systems where solution failure frequently occurs. Consequently, it also demands a high level of precision from the reference trajectory.
This letter presents a novel framework that integrates a neural network-based planner with an improved CBF-based MPC controller to ensure safety and real-time performance in robot motion control. The proposed framework serves both as a real-time local planner and as a controller.
I-A Related Work
I-A1 Path Planner
Sampling-based and search-based methods have been extensively studied, such as A*[4], PRM[5], and RRT[6]. These methods are fundamental in the field of robotics for path planning and have been widely researched and implemented in various applications. However, these methods often suffer from high computational costs and prolonged computation times, thus being only applicable to low-dimensional spaces.
Learning-based path planning methods are capable of rapidly generating high-quality paths, adapting to complex environments, and reducing reliance on prior knowledge[7][8]. For instance, neural A*[9], which is based on deep learning, offers a more efficient approach to path generation. However, since these methods can only plan paths offline, while the controllers can effectively track the paths, the prerequisite is the ability to adjust control inputs in real time. Optimization-based methods can achieve real-time path planning but often exhibit slower planning speeds[10][11].
Figure 1: Schematic representation of the motion planning and control framework integrating a neural network-based planner and a CBF-based MPC controller for obstacle avoidance.
I-A2 Optimization-Based Controller
Collision-free optimal control means finding the maximum or minimum value of the cost function under constraints. For instance, the TEB[12] algorithm achieves motion control by solving a multi-objective optimization problem. MPC has demonstrated significant potential in collision-free tasks[13].
The CBF serves as a safety boundary function for the control process, providing assurance of the robot’s safety[2][14]. CBF is particularly well-suited for collision avoidance constraints and have numerous applications in the fields of robotic manipulators [15] and vehicular systems [16]. In discrete-time systems, discrete-time control barrier function (DCBF) shows good performance in terms of safety constraints [17]. Research has been conducted on DCBF in the control and planning problems of multi-step optimization [18].
CBF-based MPC combines the advantages of both and has been increasingly used in recent years[19]. Similarly, DCBF can also be combined with MPC to achieve better results in obstacle avoidance problems [20]. The corresponding control method has been validated on both bipedal robots[21] and vehicle systems[22]. However, the above methods either oversimplify the control scenarios or have difficulty meeting the real-time requirements of control.
I-B Contributions
The contributions of this paper are as follows:
  • We propose an encoder-decoder path planning architecture which can determine the subsequent target point using historical path information and employs the Dubins curve to the target point as the reference path input to the control module.
  • Based on our planning architecture, we have improved the MPC-DCBF formulation to better accommodate nonlinear kinematic constraints and ensure its real-time performance.
  • We conducted numerical simulations and real-world experiments for the proposed framework in this letter, and performed comparative experiments with several baseline schemes.
I-C Paper Structure
The structure of this paper is organized as follows: In Section II, we present the preliminaries of MPC and CBF. In Section III, we introduce our proposed framework, along with the detailed design of the planner and controller. In Section IV, we describe our numerical simulations and real-world experiments. Finally, Section V concludes the paper.
II Preliminaries
In this section, we will introduce MPC and the optimization formulation using DCBF as the collision avoidance constraint.
II-A Model Predictive Control
The state transition equation of the robot is formulated as follows:
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t}),$ (1)
where $\mathbf{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathbf{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ denote the state and control input of the robot at time step $t$ respectively, and $f$ is continuous.
If the distance to obstacles is employed as a collision avoidance constraint, then the MPC control formulation (MPC-DC) at time $t$ is given as follows:
$\min_{\mathbf{u}_{t:t+N-1|t}}$ $\quad p(\mathbf{x}_{t+N|t})+\sum_{k=0}^{N-1}q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ (2a)
s.t. $\quad\mathbf{x}_{t+k+1|t}=f(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t}),$
$\quad\quad\text{for}\ k=0,\ldots,N-1$ (2b)
$\quad\mathbf{x}_{t+k|t}\in\mathcal{X},\,\mathbf{u}_{t+k|t}\in\mathcal{U},$
$\quad\quad\text{for}\,k=0,\ldots,N-1$ (2c)
$\quad\mathbf{x}_{t|t}=\mathbf{x}_{t},\ \mathbf{x}_{t+N|t}\in\mathcal{X}_{f}$ (2d)
$\quad g(\mathbf{x}_{t+k|t})\geq 0,\,k=0,\ldots,N-1.$ (2e)
Here, The functions $p(\cdot)$ and $q(\cdot)$ in (2a) represent the terminal cost and the cost associated with tracking the reference trajectory, respectively. Constraint (2d) indicates that the initial state of the robot is $\mathbf{x}_{t}$ and the subsequent states are predicted via the system dynamics given in (2b), and constraint (2d) also specifies the terminal constraint. Both state and input constraints are provided by (2c). The collision avoidance constraint is given by (2e), which can be defined under various Euclidean norms.
We can also regard the MPC-DC as a path planning module. The path planning problem can be formulated as planning a collision-free trajectory for the robot from the initial state $\mathbf{x}_{0}$ to the target state $\mathbf{x}_{goal}$. If we set the value of $k$ in the optimization model (2) to 0 and solve for the optimal state $\mathbf{x}$, this becomes a standard path planning problem. We will employ an encoder-decoder model to address the path planning problem, the specifics of which will be discussed in Section III-A.
II-B Control Barrier Function
If the dynamical system (1) is safe with respect to the set $S\subseteq\mathcal{X}$, then any trajectory initiated from within $S$ will remain within $S$. The set $S$ is defined as the 0-superlevel set of a continuous function $h:\mathcal{X}\rightarrow\mathbb{R}$ as:
$\mathcal{S}:=\{\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{n}:h(\mathbf{x})\geq 0\}.$ (3)
We refer to $S$ as the safety set, which encompasses all regions without obstacles. $h$ is defined as a DCBF, if satisfied
$h(f(\mathbf{x},\mathbf{u}))\geq\gamma(\mathbf{x})h(\mathbf{x}),$ (4)
$0\leq\gamma(\mathbf{x})<1,\forall\mathbf{x}\in\mathcal{S},\exists\mathbf{u}\in\mathcal{U},$
Let $\gamma_{k}:=\gamma(\mathbf{x}_{k})$. Satisfying (3) implies $h(f(\mathbf{x},\mathbf{u}))\geq\gamma_{K}h(\mathbf{x})$, that is, the lower bound of the DCBF decreases exponentially with the decay rate $\gamma_{k}$ [18].
$\mathcal{K}(\mathbf{x}):=\{\mathbf{u}\in\mathcal{U}:h(f(\mathbf{x},\mathbf{u}))-\gamma(\mathbf{x})h(\mathbf{x})\geq 0\}.$ (5)
Then, if the initial state is within the safety set $S$ and $\mathbf{u}_{k}\in\mathcal{K}(\mathbf{x}_{k})$, all subsequent states will also remain within the safety set $S$, which implies that the resulting trajectory is safe.
The formulation proposed later in [nmpc dcbf 33] introduces a slack variable $\omega$ to balance feasibility and safety as follows:
$h(f(\mathbf{x},\mathbf{u}))\geq\omega(\mathbf{x})\gamma(\mathbf{x})h(\mathbf{x}),\quad 0\leq\gamma(\mathbf{x})<1.$ (6)
The MPC control formulation based on DCBF constraints (MPC-DCBF) at time $t$ is presented as follows:
$\min_{U,\Omega}$ $\quad p(\mathbf{x}_{t+N|t})+\sum_{k=0}^{N-1}q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})+\psi(\omega_{k})$ (7a)
s.t. $\quad\mathbf{x}_{t+k+1|t}=f(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t}),$
$\quad\quad\text{for}\ k=0,\ldots,N-1$ (7b)
$\quad\mathbf{x}_{t+k|t}\in\mathcal{X},\,\mathbf{u}_{t+k|t}\in\mathcal{U},$
$\quad\quad\text{for}\,k=0,\ldots,N-1$ (7c)
$\quad h(\mathbf{x}_{t+k+1|t})\geq\omega_{k}\gamma_{k}h(\mathbf{x_{t+k|t}}),\ \omega_{k}\geq 0,$
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1,$ (7d)
$\quad\mathbf{x}_{t|t}=\mathbf{x}_{t},\ \mathbf{x}_{t+N|t}\in\mathcal{X}_{f}.$ (7e)
The optimization formulation (7) is fundamentally similar to the optimization formulation (2). In formulation (7), $U=[\mathbf{u}^{T}_{t|t},\ldots,\mathbf{u}^{T}_{t+N-1|t}]$ and $\Omega=[\omega_{0},\ldots,\omega_{N_{CBF-1}}]$ are the joint input and relaxation variables, respectively. $\psi$ is the penalty function for the relaxation variable.
Figure 2: Figure illustrating the Neural Dubins Model.
III NMPCB Motion Control Framework
In this section, we introduce the NMPCB framework, which consists mainly of two components: neural network-based path planner and the DCBF-based MPC controller.
III-A Neural Dubins Model
In this section, we describe our model formulation. As shown in Fig. 2 and Fig. 3, our model mainly consists of three components: the reference path encoder, the obstacle encoder, and the decoder.
The reference path encoder receives information from the start point $\mathbf{x}_{0}=[x_{0},y_{0},\theta_{0}]$ and the goal point $\mathbf{x}_{goal}=[x_{goal},y_{goal},\theta_{goal}]$ as input, where $x$ and $y$ represent the Cartesian coordinates, and $\theta$ denotes the heading angle. Within the model, a Dubins curve[23] is first generated from the $\mathbf{x}_{0}$ to $\mathbf{x}_{goal}$, and the coordinates of each point $P_{k}=[x_{k},y_{k},\theta_{k}]$ on the curve along are sequentially input into a single-layer LSTM module to obtain the hidden state $H$. LSTM networks exhibit superior adaptability to motion continuity and enhanced robustness to noisy or irregular path data when processing path sequences.
The obstacle encoder receives obstacle information as input. Each obstacle can be represented as $O_{i}=[x_{i},y_{i},r_{i}]$, where $x$ and $y$ represent the center, and $r$ denotes the obstacle radius. The model maps the obstacle information onto a $50\times 50$ map matrix, where matrix values are set to $1$ at grid points with obstacles and 00 at obstacle-free points, thus, the dimension of each map matrix is $(50,50,1)$. This map matrix is then fed into a hierarchical feature extraction architecture comprising two convolutional downsampling modules (CDM) and a fully connected compression module, which ultimately outputs a low-dimensional feature vector $Z$. Each convolutional downsampling module comprises individual two-dimensional convolutional layers with a kernel size of $3\times 3$. Each convolutional layer is connected with batch normalization (BN), an activation function, the rectified linear unit (ReLU), and a downsampling layer, max pooling, with filters of size $2\times 2$.
Figure 3: Figure illustrating the CDM.
Subsequently, the hidden state $H$ of the reference path encoder, the feature vector $Z$ of the obstacle encoder, the endpoint $\mathbf{x}_{goal}$, and all path points from the start point to the current point $\mathbf{x}_{0:t}$ are input into the decoder. It then outputs the future states $\hat{\mathbf{{x}}}_{t+N_{pred}}$ within a specified time horizon $N_{pred}$,
$\mathcal{D}(H,Z,\mathbf{x}_{goal},\mathbf{x}_{0:t})\rightarrow\hat{\mathbf{{x}}}_{t+N_{pred}}.$ (8)
The decoder network $\mathcal{D}$ is composed of a single-layer LSTM network followed by a fully connected layer. We use a mean-squared-error (MSE) loss between the predicted state $\hat{\mathbf{{x}}}_{t+N_{pred}}$ and the label next state $\mathbf{{x}}_{t+N_{pred}}$ during training, as follows:
$loss=\frac{1}{N}\sum_{k=0}^{N}\left\|\hat{\mathbf{x}}_{k+N_{pred}}-\mathbf{x}_{k+N_{pred}}\right\|^{2}.$ (9)
Here, $N$ denotes the total number of planning steps.
After deriving the predicted state $\hat{\mathbf{{x}}}_{t+N_{pred}}$, a Dubins curve is constructed from the current point $\mathbf{x}_{t}$ to $\hat{\mathbf{{x}}}_{t+N_{pred}}$ to serve as a reference trajectory that is transmitted to the control module.
III-B Dual DCBF Constraint
To simplify the formulation of the obstacle avoidance problem, we introduce the following assumptions. First, the robot can be abstracted into a point mass model. Second, the geometry of any obstacle can be over-approximated with a union of convex polytopes, which is Defined as a bounded polyhedron. In the l-dimensional space, the $i$-th obstacle can be described as:
$\mathcal{O}_{i}:=\{y\in\mathbb{R}^{l}:A^{\mathcal{O}_{i}}y\leq b^{\mathcal{O}_{i}}\}$ (10)
where $b^{\mathcal{O}_{i}}\in R^{s^{\mathcal{O}_{i}}}$. $s^{\mathcal{O}_{i}}$ represent the number of facets of polytopic sets for the $i$-th obstacle.
Let the state of the robot be $\mathbf{x}\in\mathbb{R}^{n}$ with its discrete-time dynamics as defined in (1) , $\mathcal{R}(\mathbf{x})$ represent the center of the robot. For $\mathbf{x}\in\mathcal{X}$, the computation of the squared minimum distance between the $i$-th obstacle $\mathcal{O}_{i}$ and the center $\mathcal{R}(\mathbf{x})$ can be be denoted by the function $h_{i}(\mathbf{x})$ formula:
$h_{i}(\mathbf{x})=$ $\min_{y^{\mathcal{O}_{i}}\in\mathbb{R}^{l},\mathbf{x}\in\mathbb{R}^{n}}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}$ (11a)
[…]
It can be noted that Equation (11) formulates a quadratic programming (QP) problem, thus representing a convex optimization problem. To ensure safety throughout the motion process, we apply the same DCBF constraint to each obstacle in relation to the robot. Thus, the safety set is defined as:
$\mathcal{S}_{i}:=\{\mathbf{x}\in\mathbb{R}^{n}:h_{i}(\mathbf{x})>0\}^{c},\text{for}\ i=1,\ldots,N_{o},$ (12)
where $(\cdot)^{c}$ denotes the closure of a set, and $N_{O}$ denotes the number of obstacles. Enforcing the DCBF constraint for each $h_{i}$ ensures that the state remains in $\mathcal{S}:=\bigcap_{i=1}^{N_{o}}S_{i}$.
We note that in the DCBF constraint (4), the computation of $h(f(\mathbf{x},\mathbf{u}))$ is required, which leads to the presence of non-differentiable implicit constraints in the optimization formulation and can consume significant computational time. Next, we introduce the dual DCBF constraints to address this issue.
For any convex optimization problem, a dual problem exists. The dual form of problem (11) is given by:
$g_{i}(\mathbf{x})=\max_{\lambda^{\mathcal{O}_{i}}}$ $\quad(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}$ (13a)
s.t. $\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1$ (13b)
$\quad\lambda^{\mathcal{O}_{i}}\geq 0$ (13c)
Here $\lambda A^{\mathcal{O}_{i}}$ is the normal vector of the separating hyperplane.
According to the Weak Duality Theorem, for all optimization problems, it holds that $g_{i}(\mathbf{x})\leq h_{i}(\mathbf{x})$ for their respective dual problems. Since (11) is a convex optimization with linear constraint and has a well-defined optimum solution in $\mathbb{R}^{+}$, the Strong Duality Theorem[24] also holds, which states that
$g_{i}(\mathbf{x})=h_{i}(\mathbf{x})$ (14)
In accordance with the Strong Duality Theorem (14), we can substitute $g_{i}(\mathbf{x})$ for the computation of $h_{i}(\mathbf{x})$, thereby circumventing the implicit dependence of $h_{i}(\mathbf{x})$ on $\mathbf{x}$. We assume $g_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}})$ represents the cost associated with any feasible solution $\lambda^{\mathcal{O}_{i}}$ of (13). Since (13) is a maximization problem, we can derive the following inequality relationship:
$\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}}):=(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\leq g_{i}(\mathbf{x})=h_{i}(\mathbf{x}).$ (15)
Then, we can transform the DCBF constraints into a more stronger form:
$(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}h_{i}(\mathbf{x}).$ (16)
This represents a stronger DCBF constraint that satisfies the (4). The substitution of the DCBF constraint with (16) requires satisfying (13b) and (13c), as follows:
$\quad(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}h_{i}(\mathbf{x}),$ (17a)
$\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1,\ \lambda^{\mathcal{O}_{i}}\geq 0.$ (17b)
According to the Strong Duality Theorem (14), $\lambda^{\mathcal{O}_{i}*}$ satisfying (10b)-(10c) such that for all $x\in\mathcal{X}$
$\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}*})=(A^{\mathcal{O}_{i}*}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}*})^{T}\lambda^{\mathcal{O}_{i}*}=h_{i}(\mathbf{x}).$ (18)
This implies that if and only if $(\mathbf{u},\lambda^{\mathcal{O}_{i}*})$ satisfy (13), for all fixed $\mathbf{x}\in\mathcal{X}$, the input $\mathbf{u}$ satisfies the DCBF constraint (5) with $h_{i}$ implicitly defined, which also implies that the set of feasible inputs does not diminish at any $\mathbf{x}\in\mathcal{X}$.
Similarly, we can apply the aforementioned method to optimization problem (11) to establish stronger DCBF constraints. Let $\mathbf{y}^{\mathcal{O}_{i}}$ be any feasible solution to (11). Since (11) is a minimization problem, we can derive the following inequality relationship:
$\bar{h}_{i}(\mathbf{x},y^{\mathcal{O}_{i}},y^{\mathcal{R}}):=\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq h_{i}(\mathbf{x}).$ (19)
Then we can enforce the stronger DCBF constrain:
$\quad(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}.$ (20)
Comparing with equation (4), we can observe the transformation of the DCBF constraints:
$h_{i}(f(\mathbf{x},\mathbf{u}))$ $\geq(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}$
$\geq\gamma_{k}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq\gamma_{k}h_{i}(\mathbf{x}).$ (21)
Integrating the analysis presented above, we define the dual DCBF constraints as follows:
$(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2},$ (22a)
$\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1,\ \lambda^{\mathcal{O}_{i}}\geq 0,$ (22b)
$A^{\mathcal{O}_{i}}y^{\mathcal{O}_{i}}\leq b^{\mathcal{O}_{i}}.$ (22c)
Introducing a slack variable $\omega$ has no effect on the aforementioned formulation.
III-C Optimization Formulation
We apply the dual DCBF constraints to the MPC algorithm as collision avoidance constraints to construct a multi-step optimization model, thereby providing enhanced assurance of safety in motion control. The optimization formulation for MPC-DUAL-DCBF (MDD) at time $t$ is presented as follows:
$\min_{U,\Omega}$ $\quad p(\mathbf{x}_{t+N|t})+\sum_{k=0}^{N-1}q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})+\psi(\omega_{k})$ (23a)
s.t. $\quad\mathbf{x}_{t|t}=\mathbf{x}_{t},\ y_{0}^{\mathcal{O}_{i}}=y_{t}^{\mathcal{O}_{i}*},$ (23b)
$\quad\mathbf{x}_{t+k|t}\in\mathcal{X},\,\mathbf{u}_{t+k|t}\in\mathcal{U},$
$\quad\quad\text{for}\,k=0,\ldots,N-1$ (23c)
$\quad\mathbf{x}_{t+k+1|t}=f(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t}),$
$\quad\quad\text{for}\ k=0,\ldots,N-1$ (23d)
$\quad(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x}_{t+k+1|t})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}_{k+1}$
$\quad\geq\omega_{k}\gamma_{k}\|y^{\mathcal{O}_{i}}_{k}-\mathcal{R}(\mathbf{x}_{t+k|t})\|_{2}^{2},$
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23e)
$\quad\|\lambda^{\mathcal{O}_{i}}_{k}A^{\mathcal{O}_{i}}\|_{2}\leq 1,\ A^{\mathcal{O}_{i}}y^{\mathcal{O}_{i}}_{k}\leq b^{\mathcal{O}_{i}},$
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23f)
$\quad\lambda^{\mathcal{O}_{i}}_{k}\geq 0,\ \omega_{k}\geq 0.$
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23g)
In the formulation, $y_{t}^{\mathcal{O}_{i}*}$ represents the optimal solution obtained by precomputing the minimum distance at time $t$ through (11). Constraint (23e) represents the dual DCBF constraint, while (23f)-(23g) denote the feasibility conditions of the dual DCBF constraint. The optimization formulation (23) illustrates only the dual DCBF constraint between obstacle $O_{i}$ and the region $R$, however, during the motion control process, the corresponding dual DCBF constraints are applied to each pair of obstacles and the robot.
If the same dual DCBF constraints are imposed at each time step within a multi-step optimization model, it would result in a substantial computational burden. To reduce this complexity, we roll out the time $k$ and replace the RHS of each DCBF constraint with $h_{i}(\mathbf{x}_{t})$, as follows:
$\quad(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x}_{t+k+1})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}_{t+k+1}\geq\omega_{k}(\prod_{j=0}^{k}\gamma_{j})h_{i}(\mathbf{x}_{t}).$ (24)
Replacing (23e) with Equation (24) can accelerate the computation. This modification affects neither the system feasibility nor its safety.
III-D Framework Overview
The framework consists of a planner and a controller. The planner utilizes Neural Dubins Model to determine the next target point and generates a Dubins curve from the current point to the target point, which serves as a reference trajectory and is passed to the control module. The control module solves equation (23) to output the optimal control commands, ensuring that the robot moves towards the goal without collisions.
We focus on the motion control problem of mobile robots, hence adopting the bicycle model as the kinematic model.The state vector of the robot $\mathbf{x}$ is defined as $[x,y,v,\theta]$, where $x$ and $y$ represent the coordinates of the rear axle center, $v$ is the velocity, and $\theta$ is the yaw angle. The control vector $\mathbf{u}$ is defined as $[a,\delta]$, where $a$ is the acceleration and $\delta$ is the steering angle of the front wheel. The kinematic equations are given by:
$x_{k+1}=x_{k}+v_{k}\cos{(\theta_{k})}\Delta t,$ (25a)
$y_{k+1}=y_{k}+v_{K}\sin{(\theta_{k})}\Delta t,$ (25b)
$v_{k+1}=v_{k}+a_{k}\Delta t,$ (25c)
$\theta_{k+1}=\theta_{k}+\frac{v_{k}\tan{\delta_{k}}}{L}\Delta t,$ (25d)
where $\Delta t$ represents the time step and $L$ denotes the wheelbase.
The cost function (23a) is composed of terminal cost, stage cost, and slack function cost, respectively as
$p(\mathbf{x}_{t+N|t})$ $=\|\mathbf{x}_{t+N|t}-\bar{\mathbf{x}}_{t+N|t}\|^{2},$ (26a)
$q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ $=\|\mathbf{x}_{t+k|t}-\bar{\mathbf{x}}_{t+k|t}\|^{2}_{Q}+\|\mathbf{u}_{t+k|t}\|^{2}$
$\quad+\|\mathbf{u}_{t+k|t}-\mathbf{u}_{t+k-1|t}\|^{2},$ (26b)
$\psi(\omega_{k})$ $=p_{\omega}(\omega_{k}-1)^{2},$ (26c)
where $\bar{\mathbf{x}}$ represents the reference trajectory. The stage cost consists of the tracking cost of the reference trajectory, the control effort cost, and the control smoothness cost.
For the point mass model of the robotic, we can ensure effective obstacle avoidance by setting a safety distance $D_{safe}$, which also alleviates the issue of decreased solution speed due to the incorporation of distant obstacles in the optimization problem.
For optimization problem (23), the geometric representation of the robot and obstacles can be altered to accommodate various scenarios. In scenarios where high solution speed is required and accurate shape information of obstacles can be obtained, an optimization formulation corresponding to a point mass robot model and convex polygon obstacles can be utilized (MDD-I). In scenarios where solution speed is not a critical factor and navigating through tight spaces is challenging, an optimization model that corresponds to a convex polygonal robot model and convex polygonal obstacles can be applied (MDD-II, [20]).
In the following section, we will conduct comparative experiments of the aforementioned algorithms across various scenarios.
IV Experiments and Results
In this section, we conduct numerical simulations of the NMPCB framework and perform ablation experiments to verify that both the planner and the controller significantly enhance the motion control problem. Furthermore, we have conducted real-world experiments with the proposed framework to validate its performance in actual environments.
(a) Dubins Curve + MPC-DCBF
(b) Dubins Curve + MDD-I
(c) Neural Dubins Model + MDD-I
(d) Neural Dubins Model + MDD-II
Figure 4: Comparison of trajectories among different algorithms in numerical simulation experiments.
IV-A Datasets Construction
Due to the scarcity of high-quality path planning datasets for deep learning in unstructured environments, we have constructed a dataset that includes information on obstacles and the robot’s motion trajectories. The generation of motion trajectories within the dataset is facilitated by the RDA [11] algorithm. We generated a substantial amount of motion trajectory data in unstructured scenarios by randomly creating obstacles and assigning start and end points within a certain range.
For ease of representation, obstacles are uniformly modeled as circles, with their information consisting of the center and radius of the circles. Trajectory information includes the coordinates of the robot’s rear axle center and the robot’s heading angle. We have excluded data where collisions occurred during motion, resulting in trajectory lengths shorter than a predetermined threshold.
We have generated a total of 12,000 instances for the training set and 3,000 instances for the test set.
IV-B Numerical Simulation
We have prepared the following four combinations of motion control algorithms for comparative experiments: (1) A Dubins curve connecting the starting point to the goal point is utilized as the planner, with the MPC-DCBF serving as the controller. (2) A Dubins curve connecting the starting point to the goal point is utilized as the planner, with the MDD-I serving as the controller. (3) The Neural Dubins Model is utilized as the planner, with the MDD-I serving as the controller. (4) The Neural Dubins Model is utilized as the planner, with the MDD-III serving as the controller. We have set the parameters for the aforementioned algorithms to $N=11$, $N_{{CBF}}=10$, and $\gamma=0.9$. The optimization problems are implemented in Python using CasADi[25] as the modeling language and solved with IPOPT [26] on Ubuntu 18.04.
We validated our proposed framework in intelligent robot simulator (ir-sim) [^1] https://github.com/hanruihua/ir_sim, which is a Python-based simulator for robotic algorithm development. We represent obstacles with black cubes, denote the area traversed by the robot with a yellow border, and illustrate the reference trajectory provided by the planner to the controller with colored lines.
In Fig. 4LABEL:sub@fig4a, the red dots indicate points where the solver failed to find a solution, demonstrating the poor performance of MPC-DCBF in complex scenarios. Fig. 4LABEL:sub@fig4b shows that MDD-I can adapt to complex environments, however, due to the Dubins curve not accounting for obstacles, the controller collided while following the reference trajectory. Fig. 4LABEL:sub@fig4c - 4LABEL:sub@fig4d illustrate that, upon employing the Neural Dubins Model, all two algorithms are capable of planning a suitable collision-free trajectory. This demonstrates that the Neural Dubins Model provides an appropriate reference path. However, there is a significant difference in computation time among the two algorithms, especially in densely obstacle-populated areas, where the computation time for MDD-II often reaches around 0.15 seconds, whereas the computation time for MDD-I does not exceed 0.1 seconds. Fig. 5 reflects the disparity in their computation times.
Figure 5: Comparison of computation cost of MDD-I and MDD-II. The horizontal axis represents the time step, and the vertical axis represents the solution time, with the unit being seconds.
To better compare the differences in success rate and average solution time among the two algorithms, we conducted comparative experiments across 50 randomly generated scenarios, and the results are presented in Table 1.
TABLE I: Performance of MDD across 50 random scenarios
Success Rate Average Solving Time Average Max Solving Time
Neural Dubins MDD-I 0.86 Four 0.172
Neural Dubins MDD-II 0.82 Four 0.851
Here, the success rate refers to the ratio of the number of scenarios where the robot successfully reaches the destination without collision to the total number of scenarios. The average solving time is the average single-step solving time in successful scenarios. The average maximum solving time refers to the average of the maximum single-step solving time in each successful scenario.
(a) Dubins Curve + MPC-DCBF
(b) Dubins Curve + MDD-I
(c) Neural Dubins Model + MDD-I
(d) Neural Dubins Model + MDD-II
Figure 6: Comparison of trajectories among different algorithms in real-world experiments. Subfigures (c) and (d) correspond to control module frequencies of 10 Hz. The red arrow indicates the location where a collision occurred.
It can be observed that MDD-I and MDD-II exhibit similar success rates. However, both average solving time and average maximum solving time for MDD-I are significantly lower than those for MDD-II.
Figure 7: The Ackermann robot.
IV-C Real-World Experiments
To validate the practical performance of the NMPCB framework in real-world scenarios, experiments were conducted on the robot depicted in Fig. 7. This robot is equipped with four wheels and employs Ackermann steering. Its sensor suite comprises a 2D LiDAR, an IMU, and wheel odometry. We utilized an NVIDIA Jetson Nano as the computing platform.
Utilizing the aforementioned hardware, we conducted comparative experiments of the four algorithms from numerical simulations in real-world settings. In Fig. 6LABEL:sub@fig6a, the MPC-DCBF algorithm faced persistent solution failures in the presence of multiple obstacles, ultimately resulting in a collision. In Fig. 6LABEL:sub@fig6b, although no solution failures occurred, collisions still took place due to the use of the Dubins curve as the reference trajectory. Comparing Fig. 6LABEL:sub@fig6c and Fig. 6LABEL:sub@fig6d, the MDD-II algorithm resulted in a collision due to the robot’s control module frequency exceeding the solution frequency of MDD-II. However, the MDD-I algorithm demonstrated superior real-time performance, successfully reaching the destination.
V CONCLUSIONS
In this paper, we propose NMPCB, a motion control framework that integrates neural networks with optimal control. This framework serves dual purposes as both a local planner and a controller. While ensuring the safety of the robot’s motion, it also demonstrates commendable real-time performance compared to other methods. For future research, we will optimize the Neural Dubins Model to enhance its capabilities in path planning. Additionally, we will focus on constructing optimization formulations of different forms to accelerate the solution process.

### v2

@@
-NMPCB: A Lightweight and Safety-Critical Motion Control Framework
+NMPCB: A Lightweight and Safety-Critical Motion Control Framework for Ackermann Mobile Robot
 Abstract
@@
 We validate the feasibility of the framework through numerical simulations and real-world experiments.
+Index Terms:
+Integrated planning and control, integrated planning and learning, collision avoidance, optimization and optimal control.
 I Introduction
 In the design of robot motion planners and controllers, safety has always been the paramount requirement, considering the rapid development of robot systems[1].
-The control method based on CBF has garnered significant attention for offering a simple yet safety-guaranteed approach to robot motion control[2].
-In particular, the control method that integrates CBF and MPC have demonstrated superior performance in safety control by predicting future states [3].
-However, this method often fails to obtain an optimal solution within a short time, especially for nonlinear kinematic systems where solution failure frequently occurs.
+The control method based on CBF has garnered significant attention for offering a simple yet safety-guaranteeing approach to robot motion control[2].
+In particular, the control method that integrates CBF and MPC has demonstrated superior performance in safety control by predicting future states [3].
+However, this method often fails to obtain an optimal solution within a short computation time, especially for nonlinear kinematic systems where solution failure frequently occurs.
 Consequently, it also demands a high level of precision from the reference trajectory.
-This letter presents a novel framework that integrates a neural network-based planner with an improved CBF-based MPC controller to ensure safety and real-time performance in robot motion control.
+This paper presents a novel framework that integrates a neural network-based planner with an improved CBF-based MPC controller to ensure safety and real-time performance in robot motion control.
 The proposed framework serves both as a real-time local planner and as a controller.
@@
 The CBF serves as a safety boundary function for the control process, providing assurance of the robot’s safety[2][14].
-CBF is particularly well-suited for collision avoidance constraints and have numerous applications in the fields of robotic manipulators [15] and vehicular systems [16].
+CBF is particularly well-suited for collision avoidance constraints and has numerous applications in the fields of robotic manipulators [15] and vehicular systems [16].
 In discrete-time systems, discrete-time control barrier function (DCBF) shows good performance in terms of safety constraints [17].
@@
 The contributions of this paper are as follows:
-  • We propose an encoder-decoder path planning architecture which can determine the subsequent target point using historical path information and employs the Dubins curve to the target point as the reference path input to the control module.
-  • Based on our planning architecture, we have improved the MPC-DCBF formulation to better accommodate nonlinear kinematic constraints and ensure its real-time performance.
-  • We conducted numerical simulations and real-world experiments for the proposed framework in this letter, and performed comparative experiments with several baseline schemes.
+  • We propose an encoder-decoder path planner that can predict a collision-free next target point using only historical trajectories and obstacle information, while significantly reducing the time overhead of the planner.
+  • We improve the MPC-DCBF formulation by constructing the dual form of obstacle avoidance constraints, which converts the implicit CBF constraints into explicit ones, thereby allowing it to better adapt to obstacle avoidance tasks and ensuring real-time performance.
+  • We conduct numerical simulations and real-world experiments for the proposed framework in this paper, and perform comparative experiments with several baseline schemes.
 I-C Paper Structure
@@
 where $\mathbf{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathbf{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$ denote the state and control input of the robot at time step $t$ respectively, and $f$ is continuous.
-If the distance to obstacles is employed as a collision avoidance constraint, then the MPC control formulation (MPC-DC) at time $t$ is given as follows:
+If the distance to obstacles is employed as a collision avoidance constraint, then the MPC control formulation at time $t$ is given as follows:
 $\min_{\mathbf{u}_{t:t+N-1|t}}$ $\quad p(\mathbf{x}_{t+N|t})+\sum_{k=0}^{N-1}q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ (2a)
@@
 $\quad\mathbf{x}_{t|t}=\mathbf{x}_{t},\ \mathbf{x}_{t+N|t}\in\mathcal{X}_{f}$ (2d)
-$\quad g(\mathbf{x}_{t+k|t})\geq 0,\,k=0,\ldots,N-1.$ (2e)
+$\quad g(\mathbf{x}_{t+k+1|t})\geq 0,\,k=0,\ldots,N-1.$ (2e)
 Here, the functions $p(\cdot)$ and $q(\cdot)$ in (2a) represent the terminal cost and the cost associated with tracking the reference trajectory, respectively.
 Constraint (2d) indicates that the initial state of the robot is $\mathbf{x}_{t}$ and the subsequent states are predicted via the system dynamics given in (2b), and constraint (2d) also specifies the terminal constraint.
-Both state and input constraints are provided by (2c).
-The collision avoidance constraint is given by (2e), which can be defined under various Euclidean norms.
-We can also regard the MPC-DC as a path planning module.
-The path planning problem can be formulated as planning a collision-free trajectory for the robot from the initial state $\mathbf{x}_{0}$ to the target state $\mathbf{x}_{goal}$.
-If we set the value of $k$ in the optimization model (2) to 0 and solve for the optimal state $\mathbf{x}$, this becomes a standard path planning problem.
+Both state and control input constraints are provided by (2c).
+The collision avoidance constraint is given by (2e), which can be defined using various Euclidean norms.
+$N$ denotes the prediction horizons.
+We can also regard the MPC as a path planning module.
+The path planning problem can be formulated as planning a collision-free trajectory for the robot from the initial state $\mathbf{x}_{0}$ to the target state $\mathbf{x}_{\text{goal}}$.
+If we set the length of the prediction horizons $N$ in the optimization model (2) to $1$ and solve for the optimal state $\mathbf{x}$, this becomes a standard path planning problem.
 We will employ an encoder-decoder model to address the path planning problem, the specifics of which will be discussed in Section III-A.
@@
 $h(f(\mathbf{x},\mathbf{u}))\geq\gamma(\mathbf{x})h(\mathbf{x}),$ (4)
-$0\leq\gamma(\mathbf{x})<1,\forall\mathbf{x}\in\mathcal{S},\exists\mathbf{u}\in\mathcal{U},$
+$0\leq\gamma(\mathbf{x})<1,\forall\mathbf{x}\in\mathcal{S},\ \exists\mathbf{u}\in\mathcal{U}.$
 Let $\gamma_{k}:=\gamma(\mathbf{x}_{k})$.
 Satisfying (4) implies $h(f(\mathbf{x},\mathbf{u}))\geq\gamma_{k}h(\mathbf{x})$, that is, the lower bound of the DCBF decreases exponentially with the decay rate $\gamma_{k}$ [18].
+We denote $\mathcal{K}(\mathbf{x})$ as
 $\mathcal{K}(\mathbf{x}):=\{\mathbf{u}\in\mathcal{U}:h(f(\mathbf{x},\mathbf{u}))-\gamma(\mathbf{x})h(\mathbf{x})\geq 0\}.$ (5)
 Then, if the initial state is within the safety set $S$ and $\mathbf{u}_{k}\in\mathcal{K}(\mathbf{x}_{k})$, all subsequent states will also remain within the safety set $S$, which implies that the resulting trajectory is safe.
-The formulation proposed later in [nmpc dcbf 33] introduces a slack variable $\omega$ to balance feasibility and safety as follows:
+The formulation proposed later in [18] introduces a slack variable $\omega$ to balance feasibility and safety as follows:
 $h(f(\mathbf{x},\mathbf{u}))\geq\omega(\mathbf{x})\gamma(\mathbf{x})h(\mathbf{x}),\quad 0\leq\gamma(\mathbf{x})<1.$ (6)
-The MPC control formulation based on DCBF constraints (MPC-DCBF) at time $t$ is presented as follows:
+The MPC control formulation based on DCBF constraint (MPC-DCBF) at time $t$ is presented as follows:
 $\min_{U,\Omega}$ $\quad p(\mathbf{x}_{t+N|t})+\sum_{k=0}^{N-1}q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})+\psi(\omega_{k})$ (7a)
@@
 $\quad h(\mathbf{x}_{t+k+1|t})\geq\omega_{k}\gamma_{k}h(\mathbf{x_{t+k|t}}),\ \omega_{k}\geq 0,$
-$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1,$ (7d)
+$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1,$ (7d)
 $\quad\mathbf{x}_{t|t}=\mathbf{x}_{t},\ \mathbf{x}_{t+N|t}\in\mathcal{X}_{f}.$ (7e)
+Figure 2: Illustration of the Neural Dubins Model.
 The optimization formulation (7) is fundamentally similar to the optimization formulation (2).
-In formulation (7), $U=[\mathbf{u}^{T}_{t|t},\ldots,\mathbf{u}^{T}_{t+N-1|t}]$ and $\Omega=[\omega_{0},\ldots,\omega_{N_{CBF-1}}]$ are the joint input and relaxation variables, respectively.
+In formulation (7), $U=[\mathbf{u}^{T}_{t|t},\ldots,\mathbf{u}^{T}_{t+N-1|t}]$ and $\Omega=[\omega_{0},\ldots,\omega_{N_{\text{CBF}-1}}]$ are the control input and relaxation variables, respectively.
 $\psi$ is the penalty function for the relaxation variable.
-Figure 2: Figure illustrating the Neural Dubins Model.
-III NMPCB Motion Control Framework
-In this section, we introduce the NMPCB framework, which consists mainly of two components: neural network-based path planner and the DCBF-based MPC controller.
-III-A Neural Dubins Model
-In this section, we describe our model formulation. As shown in Fig. 2 and Fig. 3, our model mainly consists of three components: the reference path encoder, the obstacle encoder, and the decoder.
-The reference path encoder receives information from the start point $\mathbf{x}_{0}=[x_{0},y_{0},\theta_{0}]$ and the goal point $\mathbf{x}_{goal}=[x_{goal},y_{goal},\theta_{goal}]$ as input, where $x$ and $y$ represent the Cartesian coordinates, and $\theta$ denotes the heading angle.
-Within the model, a Dubins curve[23] is first generated from the $\mathbf{x}_{0}$ to $\mathbf{x}_{goal}$, and the coordinates of each point $P_{k}=[x_{k},y_{k},\theta_{k}]$ on the curve along are sequentially input into a single-layer LSTM module to obtain the hidden state $H$.
-LSTM networks exhibit superior adaptability to motion continuity and enhanced robustness to noisy or irregular path data when processing path sequences.
-The obstacle encoder receives obstacle information as input.
-Each obstacle can be represented as $O_{i}=[x_{i},y_{i},r_{i}]$, where $x$ and $y$ represent the center, and $r$ denotes the obstacle radius.
-The model maps the obstacle information onto a $50\times 50$ map matrix, where matrix values are set to $1$ at grid points with obstacles and 00 at obstacle-free points, thus, the dimension of each map matrix is $(50,50,1)$.
-This map matrix is then fed into a hierarchical feature extraction architecture comprising two convolutional downsampling modules (CDM) and a fully connected compression module, which ultimately outputs a low-dimensional feature vector $Z$.
-Each convolutional downsampling module comprises individual two-dimensional convolutional layers with a kernel size of $3\times 3$.
-Each convolutional layer is connected with batch normalization (BN), an activation function, the rectified linear unit (ReLU), and a downsampling layer, max pooling, with filters of size $2\times 2$.
-Figure 3: Figure illustrating the CDM.
-Subsequently, the hidden state $H$ of the reference path encoder, the feature vector $Z$ of the obstacle encoder, the endpoint $\mathbf{x}_{goal}$, and all path points from the start point to the current point $\mathbf{x}_{0:t}$ are input into the decoder.
-It then outputs the future states $\hat{\mathbf{{x}}}_{t+N_{pred}}$ within a specified time horizon $N_{pred}$,
-$\mathcal{D}(H,Z,\mathbf{x}_{goal},\mathbf{x}_{0:t})\rightarrow\hat{\mathbf{{x}}}_{t+N_{pred}}.$ (8)
-The decoder network $\mathcal{D}$ is composed of a single-layer LSTM network followed by a fully connected layer.
-We use a mean-squared-error (MSE) loss between the predicted state $\hat{\mathbf{{x}}}_{t+N_{pred}}$ and the label next state $\mathbf{{x}}_{t+N_{pred}}$ during training, as follows:
-$loss=\frac{1}{N}\sum_{k=0}^{N}\left\|\hat{\mathbf{x}}_{k+N_{pred}}-\mathbf{x}_{k+N_{pred}}\right\|^{2}.$ (9)
-Here, $N$ denotes the total number of planning steps.
-After deriving the predicted state $\hat{\mathbf{{x}}}_{t+N_{pred}}$, a Dubins curve is constructed from the current point $\mathbf{x}_{t}$ to $\hat{\mathbf{{x}}}_{t+N_{pred}}$ to serve as a reference trajectory that is transmitted to the control module.
-III-B Dual DCBF Constraint
+$N_{\text{CBF}}$ denotes the safety horizon, which must satisfy $N_{\text{CBF}}\leq N$.
+II-C Obstacle Avoidance Problem
 To simplify the formulation of the obstacle avoidance problem, we introduce the following assumptions.
@@
 In the $l$-dimensional space, the $i$-th obstacle can be described as:
-$\mathcal{O}_{i}:=\{y\in\mathbb{R}^{l}:A^{\mathcal{O}_{i}}y\leq b^{\mathcal{O}_{i}}\}$ (10)
-where $b^{\mathcal{O}_{i}}\in R^{s^{\mathcal{O}_{i}}}$. $s^{\mathcal{O}_{i}}$ represent the number of facets of polytopic sets for the $i$-th obstacle.
-Let the state of the robot be $\mathbf{x}\in\mathbb{R}^{n}$ with its discrete-time dynamics as defined in (1) , $\mathcal{R}(\mathbf{x})$ represent the center of the robot.
-For $\mathbf{x}\in\mathcal{X}$, the computation of the squared minimum distance between the $i$-th obstacle $\mathcal{O}_{i}$ and the center $\mathcal{R}(\mathbf{x})$ can be be denoted by the function $h_{i}(\mathbf{x})$ formula:
+$\mathcal{O}_{i}:=\{y\in\mathbb{R}^{l}:A^{\mathcal{O}_{i}}y\leq b^{\mathcal{O}_{i}}\},$ (8)
+where $b^{\mathcal{O}_{i}}\in R^{s^{\mathcal{O}_{i}}}$. $s^{\mathcal{O}_{i}}$ represents the number of facets of polytopic sets for the $i$-th obstacle.
+Let the state of the robot be $\mathbf{x}\in\mathbb{R}^{n}$ with its discrete-time dynamics as defined in (1) , and let $\mathcal{R}(\mathbf{x})$ represent the center of the robot.
+For $\mathbf{x}\in\mathcal{X}$, the computation of the squared minimum distance between the $i$-th obstacle $\mathcal{O}_{i}$ and the center $\mathcal{R}(\mathbf{x})$ can be denoted by the function $h_{i}(\mathbf{x})$ formula:
 $h_{i}(\mathbf{x})=$ $\min_{y^{\mathcal{O}_{i}}\in\mathbb{R}^{l},\mathbf{x}\in\mathbb{R}^{n}}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}$ (9a)
-[…]
-It can be noted that Equation (11) formulates a quadratic programming (QP) problem, thus representing a convex optimization problem.
+$\quad\quad\text{s.t.}\quad\quad A^{\mathcal{O}_{i}}y^{\mathcal{O}_{i}}\leq b^{\mathcal{O}_{i}}.$ (9b)
+It can be noted that (9) formulates a quadratic programming (QP) problem, thus representing a convex optimization problem.
 To ensure safety throughout the motion process, we apply the same DCBF constraint to each obstacle in relation to the robot.
@@
 Enforcing the DCBF constraint for each $h_{i}$ ensures that the state remains in $\mathcal{S}:=\bigcap_{i=1}^{N_{o}}S_{i}$.
+III NMPCB Motion Control Framework
+In this section, we introduce the NMPCB framework, which consists mainly of two components: the neural network-based path planner and the DCBF-based MPC controller.
+III-A Neural Dubins Model
+In this section, we describe our encoder-decoder model formulation.
+As shown in Fig.
+2 and Fig.
+3, our model mainly consists of three components: the reference path encoder, the obstacle encoder, and the decoder.
+The reference path encoder receives information from the start point $\mathbf{x}_{0}=[x_{0},y_{0},\theta_{0}]$ and the goal point $\mathbf{x}_{\text{goal}}=[x_{\text{goal}},y_{\text{goal}},\theta_{\text{goal}}]$ as input, where $x$ and $y$ represent the position of the robot, and $\theta$ denotes the heading angle of the robot.
+Within the model, a Dubins curve[23] is first generated from $\mathbf{x}_{0}$ to $\mathbf{x}_{\text{goal}}$, and the coordinates of each point $P_{k}=[x_{k},y_{k},\theta_{k}]$ on the curve are sequentially input into a single-layer LSTM [24] module to obtain the hidden state $H$.
+LSTM networks exhibit superior adaptability to motion continuity and enhanced robustness to noisy or irregular path data when processing path sequences.
+The obstacle encoder receives obstacle information as input.
+Each obstacle can be represented as $O_{i}=[x_{i},y_{i},r_{i}]$, where $x$ and $y$ represent the center, and $r$ denotes the obstacle radius.
+The model maps the obstacle information onto a $50\times 50$ map matrix, where matrix values are set to $1$ at grid points with obstacles and 0 at obstacle-free points, thus, the dimension of each map matrix is $(50,50,1)$.
+This map matrix is then fed into a hierarchical feature extraction architecture comprising two convolutional downsampling modules (CDM) and a fully connected compression module, which ultimately outputs a low-dimensional feature vector $Z$.
+Each CDM comprises individual two-dimensional convolutional layers with a kernel size of $3\times 3$.
+Each convolutional layer is followed by batch normalization (BN), ReLU activation, and a $2\times 2$ max pooling layer for downsampling.
+Figure 3: Illustration of the CDM.
+Subsequently, the hidden state $H$ of the reference path encoder, the feature vector $Z$ of the obstacle encoder, the endpoint $\mathbf{x}_{\text{goal}}$, and all path points from the start point to the current point $\mathbf{x}_{0:t}$ are input into the decoder.
+It then outputs the future states $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$ within a specified time horizon $N_{\text{pred}}$,
+$\mathcal{D}(H,Z,\mathbf{x}_{\text{goal}},\mathbf{x}_{0:t})\rightarrow\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}.$ (11)
+The decoder network $\mathcal{D}$ is composed of a single-layer LSTM network followed by a fully connected layer.
+We use a mean squared error (MSE) loss between the predicted state $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$ and the label next state $\mathbf{{x}}_{t+N_{\text{pred}}}$ during training.
+The loss function $\mathcal{L}$ is defined as follows:
+$\mathcal{L}(\hat{\mathbf{x}},\mathbf{x})=\frac{1}{N}\sum_{k=0}^{N}\left\|\hat{\mathbf{x}}_{k+N_{\text{pred}}}-\mathbf{x}_{k+N_{\text{pred}}}\right\|^{2}_{2}.$ (12)
+Here, $N$ denotes the total number of planning steps.
+After deriving the predicted state $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$, a Dubins curve is constructed from the current point $\mathbf{x}_{t}$ to $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$ to serve as a reference trajectory that is transmitted to the control module.
+We refer to the above-mentioned neural network-based path planner as the Neural Dubins Model.
+III-B Dual DCBF Constraint
 We note that in the DCBF constraint (4), the computation of $h(f(\mathbf{x},\mathbf{u}))$ is required, which leads to the presence of non-differentiable implicit constraints in the optimization formulation and can consume significant computational time.
-Next, we introduce the dual DCBF constraints to address this issue.
+Next, we introduce the dual DCBF constraint to address this issue.
 For any convex optimization problem, a dual problem exists. The dual form of problem (9) is given by:
 $g_{i}(\mathbf{x})=\max_{\lambda^{\mathcal{O}_{i}}}$ $\quad(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}$ (13a)
-s.t. $\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1$ (13b)
-$\quad\lambda^{\mathcal{O}_{i}}\geq 0$ (13c)
-Here $\lambda A^{\mathcal{O}_{i}}$ is the normal vector of the separating hyperplane.
+s.t. $\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1,$ (13b)
+$\quad\lambda^{\mathcal{O}_{i}}\geq 0.$ (13c)
+Here $(A^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}$ is the normal vector of the separating hyperplane.
 According to the Weak Duality Theorem, for all optimization problems, it holds that $g_{i}(\mathbf{x})\leq h_{i}(\mathbf{x})$ for their respective dual problems.
 Since (9) is a convex optimization with linear constraint and has a well-defined optimum solution in $\mathbb{R}^{+}$, the Strong Duality Theorem[25] also holds, which states that
-$g_{i}(\mathbf{x})=h_{i}(\mathbf{x})$ (14)
+$g_{i}(\mathbf{x})=h_{i}(\mathbf{x}).$ (14)
 In accordance with the Strong Duality Theorem (14), we can substitute $g_{i}(\mathbf{x})$ for the computation of $h_{i}(\mathbf{x})$, thereby circumventing the implicit dependence of $h_{i}(\mathbf{x})$ on $\mathbf{x}$.
@@
 $\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}}):=(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\leq g_{i}(\mathbf{x})=h_{i}(\mathbf{x}).$ (15)
-Then, we can transform the DCBF constraints into a more stronger form:
+Then, we can transform the DCBF constraint into a stronger form:
 $(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}h_{i}(\mathbf{x}).$ (16)
-This represents a stronger DCBF constraint that satisfies the (4). The substitution of the DCBF constraint with (16) requires satisfying (13b) and (13c), as follows:
+This represents a stronger DCBF constraint that satisfies the DCBF constraint (4). The substitution of the DCBF constraint with (16) requires satisfying (13b) and (13c), as follows:
 $\quad(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}h_{i}(\mathbf{x}),$ (17a)
 $\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1,\ \lambda^{\mathcal{O}_{i}}\geq 0.$ (17b)
-According to the Strong Duality Theorem (14), $\lambda^{\mathcal{O}_{i}*}$ satisfying (10b)-(10c) such that for all $x\in\mathcal{X}$
-$\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}*})=(A^{\mathcal{O}_{i}*}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}*})^{T}\lambda^{\mathcal{O}_{i}*}=h_{i}(\mathbf{x}).$ (18)
+According to the Strong Duality Theorem (14), $\exists\ \lambda^{\mathcal{O}_{i}*}$ satisfying (13b) and (13c) such that for all $x\in\mathcal{X}$,
+$\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}*})=(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}*}=h_{i}(\mathbf{x}).$ (18)
 This implies that if and only if $(\mathbf{u},\lambda^{\mathcal{O}_{i}*})$ satisfy (16), for all fixed $\mathbf{x}\in\mathcal{X}$, the input $\mathbf{u}$ satisfies the DCBF constraint (4) with $h_{i}$ implicitly defined, which also implies that the set of feasible inputs does not diminish at any $\mathbf{x}\in\mathcal{X}$.
-Similarly, we can apply the aforementioned method to optimization problem (11) to establish stronger DCBF constraints.
+Similarly, we can apply the aforementioned method to optimization problem (9) to establish stronger DCBF constraint.
 Let $\mathbf{y}^{\mathcal{O}_{i}}$ be any feasible solution to (9).
 Since (9) is a minimization problem, we can derive the following inequality relationship:
-$\bar{h}_{i}(\mathbf{x},y^{\mathcal{O}_{i}},y^{\mathcal{R}}):=\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq h_{i}(\mathbf{x}).$ (19)
+$\bar{h}_{i}(\mathbf{x},y^{\mathcal{O}_{i}}):=\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq h_{i}(\mathbf{x}).$ (19)
 Then we can enforce the stronger DCBF constrain:
 $\quad(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}.$ (20)
-Comparing with equation (4), we can observe the transformation of the DCBF constraints:
+Comparing with (4), we can observe the transformation of the DCBF constraint:
 $h_{i}(f(\mathbf{x},\mathbf{u}))$ $\geq(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}$
 $\geq\gamma_{k}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq\gamma_{k}h_{i}(\mathbf{x}).$ (21)
-Integrating the analysis presented above, we define the dual DCBF constraints as follows:
+Integrating the analysis presented above, we define the dual DCBF constraint as follows:
 $(A^{\mathcal{O}_{i}}\mathcal{R}(f(\mathbf{x},\mathbf{u}))-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}\geq\gamma_{k}\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2},$ (22a)
@@
 III-C Optimization Formulation
-We apply the dual DCBF constraints to the MPC algorithm as collision avoidance constraints to construct a multi-step optimization model, thereby providing enhanced assurance of safety in motion control.
+We apply the dual DCBF constraint to the MPC algorithm as collision avoidance constraints to construct a multi-step optimization model, thereby providing enhanced assurance of safety in motion control.
 The optimization formulation for MPC-DUAL-DCBF (MDD) at time $t$ is presented as follows:
@@
 $\quad\geq\omega_{k}\gamma_{k}\|y^{\mathcal{O}_{i}}_{k}-\mathcal{R}(\mathbf{x}_{t+k|t})\|_{2}^{2},$
-$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23e)
+$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1$ (23e)
 $\quad\|\lambda^{\mathcal{O}_{i}}_{k}A^{\mathcal{O}_{i}}\|_{2}\leq 1,\ A^{\mathcal{O}_{i}}y^{\mathcal{O}_{i}}_{k}\leq b^{\mathcal{O}_{i}},$
-$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23f)
+$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1$ (23f)
 $\quad\lambda^{\mathcal{O}_{i}}_{k}\geq 0,\ \omega_{k}\geq 0.$
-$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23g)
+$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1$ (23g)
 In the formulation, $y_{t}^{\mathcal{O}_{i}*}$ represents the optimal solution obtained by precomputing the minimum distance at time $t$ through (9).
 Constraint (23e) represents the dual DCBF constraint, while (23f)-(23g) denote the feasibility conditions of the dual DCBF constraint.
-The optimization formulation (23) illustrates only the dual DCBF constraint between obstacle $O_{i}$ and the region $R$, however, during the motion control process, the corresponding dual DCBF constraints are applied to each pair of obstacles and the robot.
-If the same dual DCBF constraints are imposed at each time step within a multi-step optimization model, it would result in a substantial computational burden.
+The optimization formulation (23) illustrates only the dual DCBF constraint between obstacle $\mathcal{O}_{i}$ and the robot $\mathcal{R}$, however, during the motion control process, the corresponding dual DCBF constraint are applied to each pair of obstacles and the robot.
+If the same dual DCBF constraint are imposed at each time step within a multi-step optimization model, it would result in a substantial computational burden.
 To reduce this complexity, we roll out the time $k$ and replace the RHS of each DCBF constraint with $h_{i}(\mathbf{x}_{t})$, as follows:
 $\quad(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x}_{t+k+1})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}_{t+k+1}\geq\omega_{k}(\prod_{j=0}^{k}\gamma_{j})h_{i}(\mathbf{x}_{t}).$ (24)
-Replacing (23e) with Equation (24) can accelerate the computation. This modification affects neither the system feasibility nor its safety.
+Replacing (23e) with (24) can accelerate the computation. This modification affects neither the system feasibility nor its safety.
 III-D Framework Overview
 The framework consists of a planner and a controller.
-The planner utilizes Neural Dubins Model to determine the next target point and generates a Dubins curve from the current point to the target point, which serves as a reference trajectory and is passed to the control module.
-The control module solves equation (23) to output the optimal control commands, ensuring that the robot moves towards the goal without collisions.
-We focus on the motion control problem of mobile robots, hence adopting the bicycle model as the kinematic model.The state vector of the robot $\mathbf{x}$ is defined as $[x,y,v,\theta]$, where $x$ and $y$ represent the coordinates of the rear axle center, $v$ is the velocity, and $\theta$ is the yaw angle.
-The control vector $\mathbf{u}$ is defined as $[a,\delta]$, where $a$ is the acceleration and $\delta$ is the steering angle of the front wheel.
+The planner utilizes Neural Dubins Model to determine the next target point and generates a Dubins curve from the current point to the predicted target point, which serves as a reference trajectory and is passed to the control module.
+The control module solves optimization problem (23) to output the optimal control commands, ensuring that the robot moves towards the goal without collisions.
+We focus on the motion control problem of mobile robots, hence adopting the bicycle model as the kinematic model.
+The state vector of the robot $\mathbf{x}$ is defined as $x,y,v,\theta]$, where $x$ and $y$ represent the coordinates of the rear axle center, $v$ is the velocity, and $\theta$ is the yaw angle.
+The control vector $\mathbf{u}$ is defined as $a,\delta]$, where $a$ is the acceleration and $\delta$ is the steering angle of the front wheel.
 The kinematic equations are given by:
@@
 $v_{k+1}=v_{k}+a_{k}\Delta t,$ (25c)
-$\theta_{k+1}=\theta_{k}+\frac{v_{k}\tan{\delta_{k}}}{L}\Delta t,$ (25d)
-where $\Delta t$ represents the time step and $L$ denotes the wheelbase.
+$\theta_{k+1}=\theta_{k}+\frac{v_{k}\tan{(\delta_{k})}}{L}\Delta t,$ (25d)
+where $\Delta t$ denotes the time step and $L$ denotes the wheelbase.
 The cost function (23a) is composed of terminal cost, stage cost, and slack function cost, respectively as
 $p(\mathbf{x}_{t+N|t})$ $=\|\mathbf{x}_{t+N|t}-\bar{\mathbf{x}}_{t+N|t}\|^{2},$ (26a)
-$q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ $=\|\mathbf{x}_{t+k|t}-\bar{\mathbf{x}}_{t+k|t}\|^{2}_{Q}+\|\mathbf{u}_{t+k|t}\|^{2}$
+$q(\mathbf{x}_{t+k|t},\mathbf{u}_{t+k|t})$ $=\|\mathbf{x}_{t+k|t}-\bar{\mathbf{x}}_{t+k|t}\|^{2}+\|\mathbf{u}_{t+k|t}\|^{2}$
 $\quad+\|\mathbf{u}_{t+k|t}-\mathbf{u}_{t+k-1|t}\|^{2},$ (26b)
@@
 where $\bar{\mathbf{x}}$ represents the reference trajectory. The stage cost consists of the tracking cost of the reference trajectory, the control effort cost, and the control smoothness cost.
-For the point mass model of the robotic, we can ensure effective obstacle avoidance by setting a safety distance $D_{safe}$, which also alleviates the issue of decreased solution speed due to the incorporation of distant obstacles in the optimization problem.
+For the point mass model of the robot, we can ensure effective obstacle avoidance by setting a safety distance $D_{\text{safe}}$, which also alleviates the issue of decreased solution speed due to the incorporation of distant obstacles in the optimization problem.
 For optimization problem (23), the geometric representation of the robot and obstacles can be altered to accommodate various scenarios.
@@
 In scenarios where solution speed is not a critical factor and navigating through tight spaces is challenging, an optimization model that corresponds to a convex polygonal robot model and convex polygonal obstacles can be applied (MDD-II, [20]).
-In the following section, we will conduct comparative experiments of the aforementioned algorithms across various scenarios.
 IV Experiments and Results
-In this section, we conduct numerical simulations of the NMPCB framework and perform ablation experiments to verify that both the planner and the controller significantly enhance the motion control problem.
+In this section, we perform numerical simulations of the NMPCB framework and conduct repeated experiments across diverse scenarios to verify that the framework can significantly improve the solution speed of the motion control problem while ensuring robot safety.
 Furthermore, we have conducted real-world experiments with the proposed framework to validate its performance in actual environments.
-(a) Dubins Curve + MPC-DCBF
-(b) Dubins Curve + MDD-I
+(a) RDA Planner + MPC-DCBF
+(b) RDA Planner + MDD-I
 (c) Neural Dubins Model + MDD-I
 (d) Neural Dubins Model + MDD-II
-Figure 4: Comparison of trajectories among different algorithms in numerical simulation experiments.
-IV-A Datasets Construction
+(e) RDA Planner + MPC-DCBF
+(f) RDA Planner + MDD-I
+(g) Neural Dubins Model + MDD-I
+(h) Neural Dubins Model + MDD-II
+Figure 4: Comparison of trajectories among different algorithms in numerical simulations.
+IV-A Datasets Construction and Training Details
 Due to the scarcity of high-quality path planning datasets for deep learning in unstructured environments, we have constructed a dataset that includes information on obstacles and the robot’s motion trajectories.
-The generation of motion trajectories within the dataset is facilitated by the RDA [11] algorithm.
-We generated a substantial amount of motion trajectory data in unstructured scenarios by randomly creating obstacles and assigning start and end points within a certain range.
+The generation of motion trajectories within the dataset is facilitated by the RDA Planner [11].
+The simulation environment is configured as a 50 m $\times$ 50 m square workspace.
+Start and goal poses are stochastically sampled within bounded regions located at the square’s opposite corners.
+The obstacle set comprises four to six obstacles.
+A subset is uniformly sampled across the entirety of the 50 m $\times$ 50 m workspace, while the remaining obstacles are stochastically positioned within a narrow corridor centered on the arena’s diagonal axis.
+Obstacle dimensions are likewise randomized within a bounded interval, ensuring that neither excessive size nor diminutive scale compromises dataset fidelity or task validity.
+Subsequently, the planner is employed to generate collision-free trajectories.
+The obstacle configurations and corresponding trajectory data are then aggregated to constitute a single training instance.
 For ease of representation, obstacles are uniformly modeled as circles, with their information consisting of the center and radius of the circles.
@@
 We have generated a total of 12,000 instances for the training set and 3,000 instances for the test set.
+Training was performed on a workstation equipped with an NVIDIA RTX 4060 GPU and 16 GB of system memory.
+The learning rate was set to 0.001, and the Adam optimizer was adopted to accelerate convergence.
+A batch size of 32 was employed to balance memory utilization and computational efficiency.
+Finally, the model was trained for 200 epochs to ensure full convergence.
 IV-B Numerical Simulation
-We have prepared the following four combinations of motion control algorithms for comparative experiments: (1) A Dubins curve connecting the starting point to the goal point is utilized as the planner, with the MPC-DCBF serving as the controller.
-(2) A Dubins curve connecting the starting point to the goal point is utilized as the planner, with the MDD-I serving as the controller.
+We have prepared the following four combinations of motion control algorithms for comparative experiments: (1) The RDA is utilized as the planner, with the MPC-DCBF serving as the controller.
+(2) The RDA is utilized as the planner, with the MDD-I serving as the controller.
 (3) The Neural Dubins Model is utilized as the planner, with the MDD-I serving as the controller.
-(4) The Neural Dubins Model is utilized as the planner, with the MDD-III serving as the controller.
-We have set the parameters for the aforementioned algorithms to $N=11$, $N_{{CBF}}=10$, and $\gamma=0.9$.
+(4) The Neural Dubins Model is utilized as the planner, with the MDD-II serving as the controller.
+We have set the parameters for the aforementioned algorithms to $N=11$, $N_{{\text{CBF}}}=10$, and $\gamma=0.9$.
 The optimization problems are implemented in Python using CasADi[26] as the modeling language and solved with IPOPT [27] on Ubuntu 18.04.
+When the degree of constraint violation of a numerical solution is less than the feasibility tolerance of the solver, the solver will regard this solution as a feasible point.
+Consequently, violations of the obstacle avoidance constraint may occur, leading to collisions.
 We validated our proposed framework in intelligent robot simulator (ir-sim) [^1] https://github.com/hanruihua/ir_sim, which is a Python-based simulator for robotic algorithm development.
-We represent obstacles with black cubes, denote the area traversed by the robot with a yellow border, and illustrate the reference trajectory provided by the planner to the controller with colored lines.
+We represent obstacles with black polygons, denote the area traversed by the robot with a yellow border, and illustrate the reference trajectory provided by the planner to the controller with colored lines.
+Comparative simulations were conducted in both square scenarios and line scenarios to comprehensively evaluate the proposed framework.
+(a) Square Scenario
+(b) Line Scenario
+Figure 5: Comparison of computation cost of three algorithms. It illustrates the deviation of the computational cost of three algorithms at each time step from the $0.1$s benchmark line in Fig. 4.
+(a) RDA Planner + MDD-I
+(b) Neural Dubins Model + MDD-I
+(c) Neural Dubins Model + MDD-II
+(d) Line Scenario with 4 Obstacles
+(e) Square Scenario with 6 Obstacles
+(f) Square Scenario with 7 Obstacles
+(g) Square Scenario with 8 Obstacles
+(h) Line Scenario with 6 Obstacles
+Figure 6: Comparison of trajectories among different algorithms in real-world experiments. During the experiments, the control module frequency was set to 10 Hz.
 In Fig.
-4LABEL:sub@fig4a, the red dots indicate points where the solver failed to find a solution, demonstrating the poor performance of MPC-DCBF in complex scenarios.
-Fig.
-4LABEL:sub@fig4b shows that MDD-I can adapt to complex environments, however, due to the Dubins curve not accounting for obstacles, the controller collided while following the reference trajectory.
-Fig.
-4LABEL:sub@fig4c - 4LABEL:sub@fig4d illustrate that, upon employing the Neural Dubins Model, all two algorithms are capable of planning a suitable collision-free trajectory.
-This demonstrates that the Neural Dubins Model provides an appropriate reference path.
-However, there is a significant difference in computation time among the two algorithms, especially in densely obstacle-populated areas, where the computation time for MDD-II often reaches around 0.15 seconds, whereas the computation time for MDD-I does not exceed 0.1 seconds.
-Fig.
-5 reflects the disparity in their computation times.
-Figure 5: Comparison of computation cost of MDD-I and MDD-II. The horizontal axis represents the time step, and the vertical axis represents the solution time, with the unit being seconds.
-To better compare the differences in success rate and average solution time among the two algorithms, we conducted comparative experiments across 50 randomly generated scenarios, and the results are presented in Table 1.
-TABLE I: Performance of MDD across 50 random scenarios
-Success Rate Average Solving Time Average Max Solving Time
-Neural Dubins MDD-I 0.86 Four 0.172
-Neural Dubins MDD-II 0.82 Four 0.851
+4LABEL:sub@fig4a and Fig.
+4LABEL:sub@fig4e, the red dots denote locations where the solver fails to obtain a feasible solution, demonstrating the limited adaptability of MPC-DCBF in complex scenarios.
+Compared with Fig.
+4LABEL:sub@fig4b and Fig.
+4LABEL:sub@fig4f, the MDD-I controller demonstrates superior adaptability to complex environments, exhibiting no solver failures.
+Similarly, inspection of Fig.
+4LABEL:sub@fig4c - LABEL:sub@fig4d and Fig.
+4LABEL:sub@fig4g - LABEL:sub@fig4h reveals that, upon integrating the Neural Dubins Model, both frameworks consistently generate collision-free trajectories that satisfy all imposed constraints.
+Nevertheless, the three frameworks exhibit markedly disparate computational overheads.
+In obstacle-dense regions, the average solution times for RDA Planner with MDD-I and Neural Dubins Model with MDD-II cluster around 0.15 seconds, whereas Neural Dubins Model with MDD-I consistently solves within 0.10 seconds.
+These quantitative discrepancies are further illustrated in Fig.
+5.
+To quantitatively and visually assess the disparities in success rates and average solution times among the algorithms, we conducted comparative experiments across 50 randomly generated square scenarios and 50 randomly generated line scenarios.
+The aggregated results are presented in Table I and Table II.
 Here, the success rate refers to the ratio of the number of scenarios where the robot successfully reaches the destination without collision to the total number of scenarios.
-The average solving time is the average single-step solving time in successful scenarios.
-The average maximum solving time refers to the average of the maximum single-step solving time in each successful scenario.
-(a) Dubins Curve + MPC-DCBF
-(b) Dubins Curve + MDD-I
-(c) Neural Dubins Model + MDD-I
-(d) Neural Dubins Model + MDD-II
-Figure 6: Comparison of trajectories among different algorithms in real-world experiments.
-Subfigures (c) and (d) correspond to control module frequencies of 10 Hz.
-The red arrow indicates the location where a collision occurred.
-It can be observed that MDD-I and MDD-II exhibit similar success rates. However, both average solving time and average maximum solving time for MDD-I are significantly lower than those for MDD-II.
-Figure 7: The Ackermann robot.
+Across all successful scenarios, we separately report the single-step average runtime for both the planner and the controller, and additionally computed the mean of the maximum single-step execution times.
+TABLE I: Performance across 50 random square scenarios
+Success Rate
+Planner
+Controller
+Max
+RDA
+MPC-DCBF
+0.16 - - -
+RDA
+MDD-I
+0.88 54ms 32ms 167ms
+Neural Dubins
+MDD-I
+0.86 5ms 33ms 86ms
+Neural Dubins
+MDD-II
+0.88 5ms 84ms 181ms
+TABLE II: Performance across 50 random line scenarios
+Success Rate
+Planner
+Controller
+Max
+RDA
+MPC-DCBF
+0.12 - - -
+RDA
+MDD-I
+0.76 63ms 44ms 176ms
+Neural Dubins
+MDD-I
+0.76 5ms 48ms 92ms
+Neural Dubins
+MDD-II
+0.80 5ms 93ms 193ms
+Owing to the prohibitively low success rate of RDA Planner with MPC-DCBF, its runtime statistics lack statistical reliability and are therefore excluded from the reported results.
+The remaining three algorithms exhibit comparable success rates.
+However, the Neural Dubins Model markedly outperforms the RDA Planner in computational efficiency.
+Moreover, MDD-I yields a significantly lower single-step solution time than MDD-II.
+Collectively, the Neural Dubins Model paired with MDD-I demonstrates a pronounced advantage in overall solution speed.
 IV-C Real-World Experiments
@@
 We utilized an NVIDIA Jetson Nano as the computing platform.
-Utilizing the aforementioned hardware, we conducted comparative experiments of the four algorithms from numerical simulations in real-world settings.
-In Fig.
-6LABEL:sub@fig6a, the MPC-DCBF algorithm faced persistent solution failures in the presence of multiple obstacles, ultimately resulting in a collision.
-In Fig.
-6LABEL:sub@fig6b, although no solution failures occurred, collisions still took place due to the use of the Dubins curve as the reference trajectory.
-Comparing Fig.
-6LABEL:sub@fig6c and Fig.
-6LABEL:sub@fig6d, the MDD-II algorithm resulted in a collision due to the robot’s control module frequency exceeding the solution frequency of MDD-II.
-However, the MDD-I algorithm demonstrated superior real-time performance, successfully reaching the destination.
-V CONCLUSIONS
+Figure 7: The Ackermann mobile robot.
+Utilizing the aforementioned hardware, we carried out real-world experimental validation.
+Inspection of Fig.
+6LABEL:sub@fig6a - LABEL:sub@fig6c reveals that, because the control module’s execution frequency exceeds the algorithm’s solution frequency, neither RDA Planner with MDD-I nor Neural Dubins Model with MDD-II is able to reach the goal.
+In contrast, Neural Dubins Model with MDD-I demonstrates superior real-time performance and successfully reaches the goal.
+Fig.
+6LABEL:sub@fig6d - LABEL:sub@fig6h show that the Neural Dubins Model with MDD-I was further evaluated under a diverse set of challenging scenarios.
+In every trial the robot successfully reached the goal, thereby confirming the algorithm’s robust feasibility in complex environments.
+V Conclusion
 In this paper, we propose NMPCB, a motion control framework that integrates neural networks with optimal control.